OpenStudy (babyslapmafro):

When finding the area enclosed by the curves x=4-y^2 and x=y+2, how do you determine which is the right-most or left-most function? I have determined that the integral runs from 0 to 3.

OpenStudy (turingtest):

Draw the curves and find the intersection points.

OpenStudy (babyslapmafro):

x=-2,1

OpenStudy (turingtest):

-2=4-y^2 -> y=sqrt(6) -2=y+2 -> y=-4 so I think you messed up